# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ 5.2 Drug-target-interaction prediction
5.2.1 Conformalized selection for heterogeneous thresholds
We then study the behavior of vanilla CP and our proposed methods for selected units by running conformalized selection (Jin and Candès, 2023b, ) to find test pairs $j\in{\mathcal{I}}_{\textnormal{test}}$ whose outcomes are greater than data-dependent thresholds $c_{j}$. Here, we set $c_{j}$ as the $0.7$-th quantile of activities for training pairs with the same target sequence as the $j$-th data, while controlling the FDR below $q\in\{0.2,0.3,\dots,0.9\}$. The method of Bao et al., (2024) is not applicable to the heterogeneous thresholds here.
@@ 5.2.1 Conformalized selection for heterogeneous thresholds
Figure 5: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug-target-interaction for vanilla conformal prediction (Vanilla_CP) and JOMI (JOMI and JOMI_rand), at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for test units selected by conformalized selection at nominal FDR levels $q\in\{0.1,0.2,\dots,0.9\}$.
@@
5.2.2 Selection under budget constraint
@@
$\mathop{\text{maximize}}_{S\subseteq\{n+1,\dots,m+m\}}$ $\quad\sum_{j\in S}\big{[}\widehat{\mu}(X_{j})-c_{j}\big{]}$ (25)
subject to $\quad\sum_{j\in S}L_{j}\leq\bar{L},$
where $\bar{L}=200$ is a budget limit. Again, as the dataset does not come with drug development costs, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i})/\bar{\mu})+|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}-1+\varepsilon_{i}$, where $\widehat{\mu}(X_{i})$ is the predicted binding affinity, and $\epsilon_{i}\sim\text{Exp}(1)$, $\varepsilon_{i}\sim\text{Unif}([0,1])$ are independent random variables.
@@
6 Application to health risk prediction
Prediction machines are also widely used in healthcare for guiding clinical decision-making. These decisions may be determined by some complicated underlying processes such as clinical resource optimization (Ahmadi-Javid et al.,, 2017; Master et al.,, 2017) or preliminary uncertainty quantification (Olsson et al.,, 2022), and once made, would have substantial consequences.
In this section, we demonstrate the application of our framework to health risk prediction settings where units are picked by some complicated underlying processes. We will consider three selection rules:
  1. (1)
    Selection with constraints. Clinical decision makers may optimize a performance measure subject to certain constraints such as budget, capacity, or fairness (Castro and Petrovic,, 2012; Kemper et al.,, 2014; Gocgun and Puterman,, 2014), and send patients to different care categories. In Section 6.1, we consider a stylized example where we minimize the total predicted ICU stay subject to total cost budget.
  2. (2)
  3. (3)
    Selecting lower prediction bounds. Practitioners may also focus on units whose preliminary prediction intervals lie above a threshold (Levitskaya,, 2023). In Section 6.3, we study selected test units whose preliminary prediction intervals have lower bounds above a threshold.
Among the above, (1) is covariate-dependent, and (2) and (3) depend on some preliminary uncertainty quantification criterion. All of them are efficiently tackled by the computation tricks in Section 4.
Our experiments use the ICU-stay data in the MIMIC-IV dataset (Johnson et al.,, 2023). Data (and features in $X$) are pre-processed using Gupta et al., (2022), with the outcome $Y$ being the length of ICU stay. We use random forests in the scikit-learn Python package to train a point prediction model $\widehat{\mu}(\cdot)$ and two quantile regression models $\widehat{q}_{1-\beta/2}(\cdot)$ and $\widehat{q}_{\beta/2}(\cdot)$ using a heldout training set. Calibration and test data are randomly split with $n=3000$ and $m=2000$.
6.1 Selection with constraints
We first study the case where test units are selected by minimizing the total predicted ICU stay time subject to a budget constraint. Given the point predictor $\widehat{\mu}(\cdot)$, we solve a Knapsack problem similar to (25) using the same approximate algorithm, where we maximize the summation of $\widehat{\mu}(X_{j})$ in the selected subset, and the cost $L_{j}$ is generated in the same way as Section 5.2.2 and then taken the ceiling.
Figure 8 shows the empirical miscoverage rate, length of prediction interval, and sizes of the selection set and reference sets. While vanilla conformal prediction is over-confident, our method achieves exact coverage for selected test units. We also see a slightly positive correlation of |^$|\widehat{\mathcal{R}}_{j}|$ and $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$.
Figure 8: Empirical miscoverage rate (left), average length of prediction interval (middle), and scatter plots for $|{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})|$ versus {|^$\{|\widehat{\mathcal{R}}_{j}|\}_{j\in{\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})}$ (right), across $N=500$ independent runs of Vanilla_CP, JOMI, and JOMI_rand. The $x$-axis of left and middle plots is the nominal miscoverage level $\alpha\in\{0.1,0.2,\dots,0.9\}$.
6.2 Selecting small-sized prediction sets
We then consider the second selection rule, where we first build preliminary conformal prediction intervals $\widehat{C}_{\alpha}^{\text{pre}}(X_{n+j})$ via the score function $S(x,y)=|y-\widehat{\mu}(x)|/\widehat{\sigma}(x)$ (Lei et al.,, 2018), and both point prediction $\widehat{\mu}(\cdot)$ and conditional standard deviation estimation are trained via random forests. We then select those test units with $|\widehat{C}_{\alpha}^{\text{pre}}(X_{n+j})|\leq 5$, i.e., the upper and lower bounds of the preliminary prediction intervals are less than $5$ days apart. This mimics the ideas in (Ren et al.,, 2023; Sokol et al.,, 2024) where small-sized prediction sets are “certified” as confident.
After selection, we leverage the method in Section 4.3 to construct $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ for all selected test units $n+j$. Since $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is a superset of the exact output $\widehat{C}_{\alpha}(X_{n+j})$, we evaluate its empirical coverage to investigate whether it is over-conservative. Also, note from (23) that it is the union of three subsets; we also evaluate the number of disjoint segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$.
The miscoverage and length of prediction sets are reported in the first two plots in Figure 9. We observe that selectively certifying short prediction intervals can lead to under-coverage (orange curve), while JOMI achieves exact coverage (purple curve) by inflating the prediction sets, meaning that the superset $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is quite tight. The average number of disjoint segments in the right plot of Figure 9, which shows that $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ is always one single interval.
Figure 9: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI when test units with short preliminary prediction sets are selected. The $x$-axis are nominal miscoverage levels $\alpha\in\{0.1,0.2,\dots,0.9\}$.
6.3 Selecting high prediction sets
Finally, we study selection rule (iii) which is also based on a preliminary prediction set constructed in the same away as Section 6.2. We imagine that practitioners select a test unit $n+j$ if the upper bound of $\widehat{C}^{\text{pre}}(X_{n+j})$ is below $6$, i.e., it appears the patient will stay in ICU less than 6 days.
The miscoverage rate, length of prediction sets, and number of disjoint segments averaged over $N=500$ independent runs of JOMI (using the method in Section 4.3) and vanilla conformal prediction are summarized in Figure 10. We see that preliminary prediction sets with low upper bounds tend to under-cover for small $\alpha$, while JOMI achieves exact coverage despite that we construct a superset of $\widehat{C}_{\alpha}(X_{n+j})$. However, JOMI may produce multiple segments for large values of $\alpha$.
Figure 10: Empirical miscoverage rate (left), average length of prediction interval (middle), and average number of segments in $\widehat{C}^{\textrm{ps}}_{\alpha}(X_{n+j})$ (right), across $N=500$ independent runs of Vanilla_CP and JOMI. The nominal miscoverage levels on the $x$-axis are $\alpha\in\{0.1,0.2,\dots,0.9\}$.
Appendix A Additional experiment results
This section collects additional experimental results that are omitted in the main text. Appendix A.1 shows results for DPP with selection rules (ii) and (iii) with the APS score, omitted in Section 5.1. Appendix A.2 shows results for DPP with selection rules (i)-(iii) with the binary score. Appendix A.3 shows results for DTI with selection rules (ii) and (iii) that are omitted in Section 5.2.
A.1 Additional results for APS score in drug property prediction
In this part, we present results for DPP when test units are selected via (ii) top-K in both calibration and test data (Appendix A.1.1), and (iii) only calibration data as reference (Appendix A.1.2).
A.1.1 Top-K selection in mixed sample
We use the same values of $(\alpha,K)$ and the same scheme of repeated experiments as in Section 5.1.1, with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$. Figure 11 shows that vanilla conformal prediction incurs a large coverage gap for selected units, while JOMI and its randomized version both achieve valid selection-conditional coverage. The coverage gap due to discretization is eliminated again by randomization. Randomization also reduces the prediction set sizes on the right panel, but again, with $\alpha=0.8$, we see the concerning issue that it suffices for vanilla CP to construct empty prediction sets for selected units to achive marginal coverage.
A.1.2 Calibration-referenced selection
Now we select test units whose predicted affinities must be greater (or smaller) than the $K$-th largest (or smallest) predictions in the calibration set. We use the choices of $(\alpha,K)$ and the scheme of repeated experiments as in Section 5.1.1 with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$.
A.2 Additional results for binary score in drug property prediction
In Figures 1314, and 15, we show the empirical selection-conditional coverage and prediction set sizes for selection rules (i)-(iii) when the prediction sets are constructed with the binary score $V(x,y)=y(1-\widehat{\mu}(x))+(1-y)\widehat{\mu}(x)$. We observe that while vanilla CP is either over-confident or under-confident, our proposed methods achieve valid coverage across all configurations. The coverage gap of the non-randomized JOMI is smaller with this specific score function compared with the APS score.
Figure 15: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug property prediction with the binary score, for test units that are (iii) greater than the $K$-th highest (or smaller than the $K$-th lowest) predicted affinites in the calibration data. Details are as Figure 12.
A.3 Additional results for drug-target-interaction prediction
@@
In this section, we collect results for DTI with (1) Covariate dependent top-K selection. We present results under selection rule (1-i) in Appendix A.3.1, (1-ii) in Appendix A.3.2 and (1-iii) in Appendix A.3.3.
A.3.1 Top-K selection for fixed K
@@
We consider the case where test units are selected via (i) top-K (or bottom-K) among predicted affinities in the test set. We apply the vanilla conformal prediction and our proposed methods with $\mathfrak{S}=2^{{\mathcal{I}}_{\textnormal{test}}}$, so that selection-conditional coverage implies FCR control. Results under selection rules (ii) and (iii) demonstrate similar patterns, which we defer to Appendix A.3.
@@
Figure 16 shows $\widehat{\textrm{Miscov}}$ (the left panel) and the average length of the prediction intervals (the right panel) for $K\in\{20,100,1000,2000,5000,10000\}$ and nominal miscoverage level $\alpha\in\{0.1,0.8\}$. We see that vanilla conformal prediction is over-confident for pairs with the highest predicted affinites, while being under-confident for the lowest. In contrast, our proposed methods achieve near-exact coverage under different nominal levels. By comparing the solid and dashed lines in the right panel, we observe that in this problem, the intrinstic uncertainty in higher-predicted-affinity pairs is larger, so a longer prediction interval is needed to achieve certain coverage.
@@
Figure 16: Empirical miscoverage (left) and prediction interval length (right) in drug-target-interaction prediction, for vanilla conformal prediction (Vanilla_CP), JOMI (JOMI), and randomized JOMI applied to test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units. The red dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$.
A.3.2 Top-K selection in mixed sample
@@
Figure 17 shows the miscoverage and average length of prediction intervals when (ii) the focal test units are those whose predicted affinities are among the $K$ highest or lowest in both calibration and test data. We see that vanilla CP is over-confident for highest-prediction units, while under-confident for lowest-prediction units. In contrast, our proposed methods always yield valid coverage.
@@
A.3.3 Calibration-referenced selection
@@
Figure 18 shows the results with (iii) calibration-referenced selection, i.e., the selected test units are those whose predicted affinities are greater than (or smaller than) the $K$-th highest (or lowest) in the calibration data. Similar to rules (i) and (ii), vanilla CP is not calibrated for these focal units, while our methods yield near-exact coverage across the whole spectrum of $K$.
@@
Figure 18: Empirical FCR in drug-target interaction prediction, for selected test units whose $\widehat{\mu}(X_{n+j})$ surpass the highest $K$ predicted affinities in the calibration data (or smaller than the lowest). Details are otherwise the same as Figure 16.
Appendix B Technical proofs
B.1 Proof of Proposition 1
[…]
@@ B.1 Proof of Proposition 1
$\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha}(X_{n+j}){\,|\,}j\in\widehat{\mathcal{S}})$ $=\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha}(X_{n+j})\,\big{|}\,j\in\widehat{\mathcal{S}},\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}\big{)}\cdot\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})$
@@
$\geq(1-\alpha)\cdot\sum_{\ell\in{\mathcal{L}}}\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})=1-\alpha,$
@@
where we use the fact that $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint and their union is the full set. Thus, it satisfies the weak selection-conditional coverage (2).
B.2 Proof of Proposition 2

### v3

@@ C.4 Application to drug-target-interaction prediction
C.4.1 Conformalized selection for heterogeneous $c_{n+j}$’s
We then study the behavior of vanilla CP and our proposed methods for selected units by running conformalized selection (Jin and Candès, 2023b, ) to find test pairs $j\in[m]$ whose outcomes are greater than data-dependent thresholds $c_{n+j}$. Here, we set $c_{n+j}$ as the $0.7$-th quantile of activities for training pairs with the same target sequence as the $j$-th data, while controlling the FDR below $q\in\{0.2,0.3,\dots,0.9\}$. The method of Bao et al., (2024) is not applicable to the heterogeneous thresholds here.
@@ C.4.1 Conformalized selection for heterogeneous $c_{n+j}$’s
Figure 17: Empirical selection-conditional miscoverage (left) and prediction set size (right) in drug-target-interaction for vanilla conformal prediction (Vanilla_CP) and JOMI (JOMI and JOMI_rand), at nominal miscoverage $\alpha\in\{0.1,0.8\}$ for test units selected by conformalized selection at nominal FDR levels $q\in\{0.2,\dots,0.9\}$.
@@
C.4.2 Selection under budget constraint
@@
$\mathop{\text{maximize}}_{S\subseteq[m]}$ $\quad\sum_{j\in S}\big{[}\widehat{\mu}(X_{n+j})-c_{n+j}\big{]}$ (29)
subject to $\quad\sum_{j\in S}L_{n+j}\leq\bar{L},$
where $\bar{L}=200$ is a budget limit. Again, as the dataset does not come with drug development costs, we generate $L_{i}=\exp(3\widehat{\mu}(X_{i})/\bar{\mu})+|\sin(\widehat{\mu}(X_{i}))|+\epsilon_{i}-1+\varepsilon_{i}$, where $\bar{\mu}=\max_{i\in{\mathcal{D}}_{\textnormal{train}}}|\widehat{\mu}(X_{i})|$, and $\epsilon_{i}\sim\text{Exp}(1)$, $\varepsilon_{i}\sim\text{Unif}([0,1])$ are independent random variables.
@@
The empirical FCR in the two tasks is reported in Supplementary Section C.6 showing similar messages as in drug property prediction tasks. While JOMI consistently controls the empirical FCR, it can in general be lower than the selection-conditional coverage when the selection set may be empty.
C.5 Additional results for drug-target-interaction prediction
@@
In this section, we collect results for DTI with (1) covariate dependent top-K selection. We present results under selection rule (1-i) in Supplementary Section C.5.1, (1-ii) in Supplementary Section C.5.2 and (1-iii) in Supplementary Section C.5.3.
C.5.1 Top-K selection for fixed K
@@
We consider the case where test units are selected via (i) top-K (or bottom-K) among predicted affinities in the test set. We apply the vanilla conformal prediction and our proposed methods with $\mathfrak{S}=2^{[m]}$, so that selection-conditional coverage implies FCR control. Results under selection rules (ii) and (iii) demonstrate similar patterns, which are in Supplementary Section C.5.
@@
Figure 20 shows $\widehat{\textrm{Miscov}}$ (the left panel) and the average length of the prediction intervals (the right panel) for $K\in\{20,100,1000,2000,5000,10000\}$ and nominal miscoverage level $\alpha\in\{0.1,0.8\}$. We see that vanilla conformal prediction is over-confident for pairs with the highest predicted affinities, while being under-confident for the lowest. In contrast, our proposed methods achieve near-exact coverage under different nominal levels. By comparing the solid and dashed lines in the right panel, we observe that in this problem, the intrinsic uncertainty in higher-predicted-affinity pairs is larger, so a longer prediction interval is needed to achieve certain coverage.
Finally, similar to DPP, BY is overly conservative despite valid coverage. The right panel reveals another drawback of BY: it doesn not adapt to the nature of selection. More specifically, by definition, it only depends on the $\alpha|\widehat{\mathcal{S}}|/m$-th quantile of calibration scores. As such, it only relies on the size of the selection set, neglecting the nature of the selection (i.e., whether we select highest predictions or lowerst predictions). In both cases, it returns a prediction set of the form $[\widehat{\mu}(X_{n+j})-\widehat{\eta},\widehat{\mu}(X_{n+j})+\widehat{\eta}]$ where $\widehat{\eta}$ remains the same between two ranking options. In contrast, by finding “similar” calibration data in the reference set, JOMI provides more precise and targeted uncertainty quantification.
@@
Figure 20: Empirical miscoverage (left) and prediction interval length (right) in drug-target-interaction prediction, for vanilla conformal prediction (Vanilla_CP), BY (BY), JOMI (JOMI), and randomized JOMI applied to test units whose $\widehat{\mu}(X_{n+j})$ are top-$K$ (highest, solid line) or bottom-$K$ (lowest, dashed line) among test units. The red dashed line is the nominal miscoverage level $\alpha\in\{0.1,0.8\}$.
C.5.2 Top-K selection in mixed sample
@@
Figure 21 shows the miscoverage and average length of prediction intervals when (ii) the focal test units are those whose predicted affinities are among the $K$ highest or lowest in both calibration and test data. We see that vanilla CP is over-confident for highest-prediction units, while under-confident for lowest-prediction units; BY is overly conservative in both cases. In contrast, our proposed methods always yield valid coverage.
@@
C.5.3 Calibration-referenced selection
@@
Figure 22 shows the results with (iii) calibration-referenced selection, i.e., the selected test units are those whose predicted affinities are greater than (or smaller than) the $K$-th highest (or lowest) in the calibration data. Similar to rules (i) and (ii), vanilla CP is not calibrated for these focal units, while our methods yield near-exact coverage across the whole spectrum of $K$. As usual, BY is overly conservative.
@@
Figure 22: Empirical miscoverage (left) and prediction interval length (right) in drug-target interaction prediction, for selected test units whose $\widehat{\mu}(X_{n+j})$ surpass the highest $K$ predicted affinities in the calibration data (or smaller than the lowest). Details are otherwise the same as Figure 20.
C.6 Additional results for empirical FCR in drug target interaction
In this part, we collect results for the empirical FCR in drug-target interaction tasks in Section C.4.
Figure 23 plots the empirical FCR when units are selected by running conformalized selection. We observe lower empirical FCR than selection-conditional coverage, demonstrating limitations of FCR as an error metric similar to the message in Section C.3.
Figure 24 plots the empirical FCR when units are selected by constrained optimization. We observe tight empirical FCR below the nominal level, and in these experiments the selection set is always nonempty.
Figure 23: Empirical FCR in drug-target interaction prediction when units are selected by conformal selection with heterogeneous $c_{n+j}$. The $x$-axis is the nominal FDR level in the selection program.
Figure 24: Empirical FCR in drug-target interaction prediction when units are selected by constrained optimization. The $x$-axis is the nominal miscoverage level $\alpha\in\{0.1,\dots,0.9\}$.
C.7 Additional results for empirical FCR in health risk prediction
Figure 25 presents the empirical FCR for three tasks in health risk prediction. In these examples, the selection set sizes are stable and nonempty, hence the empirical FCR is also tightly controlled.
Figure 25: Empirical FCR in health risk prediction when units are selected by constrained optimization (left), length of preliminary prediction interval (middle), and lower bound of preliminary prediction interval (right). The $x$-axis is the nominal miscoverage level $\alpha\in\{0.1,\dots,0.9\}$.
Appendix D Technical proofs
D.1 Proof of Proposition 1
Suppose $\widehat{C}_{\alpha,n+j}^{(\ell)}$ satisfies $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\geq 1-\alpha$ for a set of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and we let $\widehat{C}_{\alpha,n+j}=\widehat{C}_{\alpha,n+j}^{(\ell)}$ whenever $j\in\widehat{\mathcal{S}}$ and $\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}$. This is well-defined because $\mathfrak{S}_{\ell}$’s are disjoint. Then
@@ D.1 Proof of Proposition 1
$\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in\widehat{\mathcal{S}})$ =$\stackrel{{\scriptstyle\textrm{(a)}}}{{=}}\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)},\widehat{{\mathcal{S}}}\in\mathfrak{S}_{\ell}\,\big{|}\,j\in\widehat{\mathcal{S}}\big{)}$
$=\sum_{\ell\in{\mathcal{L}}}\mathbb{P}\big{(}Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}\,\big{|}\,j\in\widehat{\mathcal{S}},\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}\big{)}\cdot\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})$
@@
$\stackrel{{\scriptstyle\textrm{(b)}}}{{\geq}}(1-\alpha)\cdot\sum_{\ell\in{\mathcal{L}}}\mathbb{P}(\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}{\,|\,}j\in\widehat{{\mathcal{S}}})=1-\alpha,$
@@
where in the first step we use the fact that $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ are disjoint and their union is the full set, and the second step follows from the strong selection-conditional coverage. By definition, it satisfies the weak selection-conditional coverage (2).
D.2 Proof of Proposition 2